A 0.200 M solution of Nacio prepared by dissolving Nacio in water. A 55.0 mL sample of this solution is is M HCl. Ka for HCio x 10 5. Calculate the pH of the solution titrated with is 3.5 at each of the following points of the titration: a. Prior to the addition of any H pH b. Halfway to the equivalence point pH c. At the equivalence point pH d. After 6.00 mL ofHCl has been added beyond the equivalence point pH
Answer #1

NaCLO + HCl _----------> HClO + NaCl

55x 0.2 0 0 0 initial mmoles

1) before tittration

pH of a salt of weak acid and strong base is given by

pH = 1/2 [pKw + pKa + log C]

= 1/2[ 14 +(8-log3.5) + log 0.2]

= 10.3785

2) Half equivalence point

[NaClO] = [HclO] and the ph of the buffer is given by Hendersen equation

pH = pKa + log [conjugate base]/[acid]

= pKa

= 7.456

3) At the equivalence

0 0 11 11 after mmoles

0 0 11/110 concentrations

The solution has only weak acid , whose pH is calculated as

pH = 1./2 [pKa -logC]

=1/2[7.456 - log 0.1]

= 3.228

4) exces Hcl added

0 1.2 11 11 mmoles after reaction

the solution has a strong acid and a weak acid. Due to common ion effect all [H+] is from strong acid only .

Thus [H+] = 1.2/116 =0.0103 M

pH = - log 0.0103

= 1.985